OpenStudy (anonymous):

Write this equation in slope-intercept form,and then find the slope and y-intercept. 5y+3x=10

OpenStudy (anonymous):

To get it in slope intercept form y=mx+b Get the y by itself on the left side of the equation

OpenStudy (anonymous):

So I divide 5y by 5 and 10 by 5 right? Then I'm left with y+3x=10 I think?

OpenStudy (anonymous):

Well yeah, but i would move the x to the right first by subtracting it. 5y=-3x+10 then divide 5y by 5 and the whole right side by 5

OpenStudy (anonymous):

And you get y=-3/5+2 Now can you see what the slope and y intercept is?

OpenStudy (anonymous):

No.Wait,You lost me., Where did you get -3 from? I subtracted 3x by 3 and 10 by 3.Which leaves me with 5y=7x right?

OpenStudy (anonymous):

Oh no you're just moving things to the right, The x is being added to the left so you subtract it from the right side, which eleminates it from the left. 5y=-3x+10 Then you do the dividing by 5

OpenStudy (anonymous):

So how did you get the -3 again?

OpenStudy (anonymous):

This is so confusing to me...>.<

OpenStudy (anonymous):

You're subtracting 3x from both sides....its like this: 5y+3x=10 -3x -3x When you subtract 3x from the left side it equals 0, so that just goes away. when you subtract it from the right you have 5y=-3x+10

OpenStudy (anonymous):

Oh..Okay.Thanks a ton!

OpenStudy (anonymous):

Then you divide by 5 and get y=-3/5x+2 and thats your answer

OpenStudy (anonymous):

I got that part.;P

OpenStudy (anonymous):

And you got the slope and y intercept?

OpenStudy (anonymous):

I don't know yet..Still working on it..

OpenStudy (anonymous):

Well its in slope intercept form y=mx+b where m=slope and b=y intercept So just look at the equation and you can get them

OpenStudy (anonymous):

The x-intercept is -3/5 and the y-intercept is 2 right?

OpenStudy (anonymous):

almost, the slope is -3/5 and the y intercept is 2.

OpenStudy (anonymous):

Oh. Thanks again!

OpenStudy (anonymous):

You're welcome